# Complete non-ambiguous trees and associated permutations: new enumerative results

### v2

@@ Abstract.
We study a link between complete non-ambiguous trees (CNATs) and permutations exhibited by Daniel Chen and Sebastian Ohlig in recent work. In this, they associate a certain permutation to the leaves of a CNAT, and show that the number of $n$-permutations that are associated with exactly one CNAT is $2^{n-2}$. We connect this to work by the first author and co-authors linking complete non-ambiguous trees and the Tutte polynomial of the associated permutation graph. This allows us to prove a number of conjectures by Chen and Ohlig on the number of $n$-permutations that are associated with exactly $k$ CNATs for various $k>1$, via various bijective correspondences between such permutations. We also exhibit a new bijection between $(n-1)$-permutations and CNATs whose permutation is the decreasing permutation $n(n-1)\cdots 1$. This bijection maps the left-to-right minima of the permutation to dots on the top row of the corresponding CNAT, and descents of the permutation to empty rows of the CNAT.
Keywords: complete non-ambiguous trees, permutations, permutation graphs, Tutte polynomial, acyclic orientations.
@@ 1. Introduction
We can associate a permutation to a CNAT by keeping only its leaf dots (see Section 2.4). This link between CNATs and permutations was first noted, to the best of our knowledge, in [13] (see Section 4 in that paper for more details). Chen and Ohlig [11] initiated the first in-depth combinatorial study of this relationship. In particular, they characterised the set of permutations $\pi$ which are associated to a unique CNAT. More generally, they looked into the number of $n$-permutations with $k$ associated CNATs, for $k\geq 1$, and provided a number of conjectures on the enumeration of such permutations. They also studied in detail so-called upper-diagonal CNATs, which are CNATs where the associated permutation is the decreasing permutation $n(n-1)\cdots 1$, using these to prove a conjecture of Laborde-Zubieta on occupied corners of tree-like tableaux [20]. Finally, they introduced a natural statistic, determinant, on CNATs, and conjectured that CNATs with determinant $+1$ and those with determinant $-1$ are equinumerous. This conjecture was solved very recently by Aval [2] through a bijective approach.
@@
In this paper, we will further develop the study initiated in [11], using results from [13]. In that work it was shown that the number of CNATs associated with a given permutation $\pi$ is equal to the number of minimal recurrent configurations for the Abelian sandpile model on the permutation graph $G_{\pi}$ of $\pi$. Minimal recurrent configurations are in natural bijection with certain acyclic orientations of the graph, and are counted by a specification of the ubiquitous Tutte polynomial (see Proposition 2.2 and Theorem 2.12 in this paper). This enumerative result allows us to solve a number of conjectures in [11], via various bijective correspondences between permutations with a given number of CNATs.
@@
We also add to the study of upper-diagonal CNATs by providing a new bijection between upper-diagonal CNATs of size $n$ and permutations of length $n-1$ (see Theorem 3.5). This new bijection has the following added benefits. Firstly, it is direct, whereas the bijection in [11] used intermediate objects called tiered trees (see [12]). Secondly, it maps certain statistics of the upper-diagonal CNATs such as the number of top-row dots, or the number of empty rows, to well-known statistics of the corresponding permutation.
@@
Our paper is organised as follows. In Section 2, we recall some necessary definitions of, and notation on, the combinatorial objects that will be considered in this paper. These include graphs, Tutte polynomials, permutations, and CNATs with associated permutations. We also show a number of useful preliminary results that will be useful in the remainder of the paper. At the end of the section we provide the key connection between CNATs and the Tutte polynomial (Theorem 2.12). In Section 3, we focus on one specific family of CNATs, called upper-diagonal CNATs. We introduce a concept of labelled CNAT, and describe a bijection between labelled CNATs and permutations of the label set (Theorem 3.5), which preserves certain statistics of both objects. This bijection then specialises to a bijection between upper-diagonal CNATs of size $n$ and permutations of length $n-1$. The bijection relies on two operations on labelled CNATs, called top-row decomposition and top-row deletion.
@@
In Section 4, we focus on counting permutations according to their number of associated CNATs. Specifically, for $k,n\geq 1$ we are interested in the set $B(n,k)$ of permutations of length $n$ which are associated with exactly $k$ CNATs. The study of these sets was initiated in [11], and we continue that work here by proving a number of conjectures left by the authors. We begin in Section 4.1 by giving characterisations of the sets $B(n,1)$ and $B(n,2)$ in terms of the so-called quadrant condition (Definition 4.1, Propositions 4.2 and 4.10). This allows us to describe a simple bijection between the product set $\{2,\cdots,n\}\times B(n,1)$ and $B(n+1,2)$ (Theorem 4.3), and deduce the enumerative formula for the latter (Corollary 4.5). In Section 4.2 we use permutation patterns to establish a bijection between the sets $B(n,2)$ and $B(n+1,3)$ (Theorem 4.16). We then show in Section 4.3 that the set $B(n,5)$ is always empty (Theorem 4.18). Finally, in Section 4.4 we consider the maximal value of $k$ such that $B(n,k)$ is non-empty. We show (Theorem 4.24) that this is achieved for $k=(n-1)!$, and that the unique permutation in $B(n,(n-1)!)$ is precisely the decreasing permutation $n(n-1)\cdots 1$ whose CNATs were studied in detail in Section 3. We conclude our paper in Section 5 with a brief summary of our results as well as some open problems and directions for future research.
@@ 2.1. Graphs and orientations
A cycle in the graph $G$ is a sequence of vertices $v_{0},v_{1},\cdots,v_{n}=v_{0}$ (for some $n\geq 3$) such that for every $i>0$, $(v_{i-1},v_{i})$ is an edge of $G$, and the vertices $v_{0},v_{1},\cdots,v_{n-1}$ are all distinct. The integer $n$ is the length of the cycle, and we will say that $G$ contains a cycle of length $n$ (or $n$-cycle for short) if there exists such a cycle in $G$. The $n$-cycle (graph) $C_{n}$ is the graph consisting of a single cycle of length $n$, with no other vertices or edges. If $G$ contains a cycle $v_{0},v_{1},\cdots,v_{n}=v_{0}$ such that $G\left[\{v_{0},\cdots,v_{n-1}\}\right]$ is the $n$-cycle $C_{n}$, we say that $G$ induces an $n$-cycle. A tree is a (connected) graph containing no cycle of any length. A spanning tree of a graph $G$ is a spanning subgraph which is a tree. The $n$-path (graph) is the graph with vertex set $V(G)=\{v_{1},\cdots,v_{n}\}$ and edge set $E(G)=\{(v_{i},v_{i+1})\}_{1\leq i\leq n-1}$ (equivalently the $n$-cycle with any single edge removed).
@@
An orientation is acyclic if it contains no directed cycle, i.e. there is no sequence $v_{0},\cdots,v_{n-1}$ of vertices such that $v_{0}\xrightarrow{\mathcal{O}}v_{1}\xrightarrow{\mathcal{O}}\cdots\xrightarrow{\mathcal{O}}v_{n-1}\xrightarrow{\mathcal{O}}v_{0}$. It is straightforward to check that an acyclic orientation must have at least one source and at least one sink. For $s\in V(G)$, we see that an acyclic orientation $\mathcal{O}$ is $s$-rooted if $s$ is the unique sink of $\mathcal{O}$.
2.2. Tutte polynomials
A key result that we will use in this paper links CNATs (see Section 2.4) to the Tutte polynomial of a graph. The Tutte polynomial was formally introduced in [24] in the author’s study graph colouring problems, although it is also implicit in earlier related work by Whitney [26]. It is a universal (in some sense, see [15, Section 2.4]) graph invariant, which appears in a number of contexts. For a detailed introduction to the Tutte polynomial and its properties, we refer readers to the book [14], in particular Chapters 2 and 3. There are three standard definitions of the Tutte polynomial. In this work, we use the definition in terms of deletion-contraction operations.
Given a graph $G$ and an edge $e=(v,w)\in E(G)$, we denote by $G\setminus e$ the graph $G$ with edge $e$ deleted. That is, $G\setminus e:=\left(V(G),E(G)\setminus\{e\}\right)$. Similarly, $G\cdot e$ denotes the graph $G$ with edge $e$ contracted. This is the graph $G$ in which the two end-points $v$ and $w$ are replaced by a single vertex $vw$, and edges incident to either $v$ or $w$ are made incident to $vw$ instead (with multiplicity preserved).
Definition 2.1.
Given a graph $G$ and an edge $e\in E(G)$, the Tutte polynomial of $G$, denoted $T_{G}$, is a bivariate polynomial defined recursively as follows.
(1) $T_{G}(x,y)=\left\{\begin{array}[]{ll}T_{G\setminus e}(x,y)+T_{G\cdot e}(x,y)&\text{if }e\text{ is neither a bridge nor a loop},\\ x\cdot T_{G\cdot e}(x,y)&\text{if }e\text{ is a bridge},\\ y\cdot T_{G\setminus e}(x,y)&\text{if }e\text{ is a loop},\\ 1&\text{if }G\text{ consists of a single vertex (and no edges)}.\\ \end{array}\right.$
In particular, the definition of $T_{G}$ does not depend on the choice of edge $e$.
We will in fact only be interested in a specific evaluation of the Tutte polynomial at $x=1$ and $y=0$. To simplify notation, we define $t_{G}:=T_{G}(1,0)$ for a graph $G$. The parameter $t_{G}$ enumerates a number of structures or quantities on the graph $G$, such as spanning trees with no external activity, minimal recurrent configurations of the Abelian sandpile model, or sink-rooted orientations. We use the latter here (see e.g. [6] for a bijective proof), and refer to $t_{G}$ as the acyclic orientation number of the graph $G$.
@@ 2.2. Tutte polynomials
Let $G$ be a graph, and $s\in V(G)$ a fixed vertex of $G$. Then $t_{G}$ is equal to the number of $s$-rooted acyclic orientations of $G$. In particular, this number does not depend on the choice of sink vertex $s$.
We now state a number of useful structural results on how the Tutte polynomial behaves with respect to certain graph constructions. We begin by characterising graphs whose acyclic orientation number is 00 or $1$.
@@
Let $G$ be a graph. Then we have $t_{G}=0$ if, and only if, $G$ contains a loop. In particular, if $G$ is simple, we have $t_{G}\geq 1$.
@@
If $G$ contains a loop, then applying the loop case of Equation (1) and evaluating at $y=0$ immediately implies that $t_{G}=0$. Conversely, if $G$ is loop-free, fix some arbitrary total order $v_{1},\cdots,v_{n}$ of the vertex set $V(G)$. Define an orientation $\mathcal{O}$ of $G$ by $v_{i}\xrightarrow{\mathcal{O}}v_{j}$ if, and only if, $i<j$. Since $G$ is loop-free, this does indeed assign a direction to each edge of $G$, and by construction $\mathcal{O}$ is a $v_{n}$-rooted acyclic orientation. Proposition 2.2 then implies that $t_{G}\geq 1$, as desired. ∎
@@
Let $G$ be a graph. Then we have $t_{G}=1$ if, and only if, all cycles in $G$ have length $2$. In particular, if $G$ is simple, then $t_{G}=1$ if, and only if, $G$ is a tree.
@@
Suppose that $G$ is such that all its cycles have length $2$, and let $e=(v,w)$ be a multi-edge corresponding to such a cycle (i.e. $e$ has multiplicity at least $2$ in $G$). By definition, we have $t_{G}=t_{G\setminus e}+t_{G\cdot e}$. But contracting $e$ in $G$ yields a loop at the new vertex $vw$, since $e$ was a multi-edge in $G$. Therefore $t_{G\cdot e}=0$ by Lemma 2.3, which implies that $t_{G}=t_{G\setminus e}$. Repeating this process until all mutli-edges have been removed, we are left with a tree $G^{\prime}$ and $t_{G}=t_{G^{\prime}}$. But now all edges in $G^{\prime}$ are bridges, and we may contract them one-by-one until we reach a graph consisting of a single vertex. Applying Equation (1) throughout this process (and evaluating at $x=1$ immediately yields $t_{G^{\prime}}=1$, as desired.
Conversely, suppose that $G$ contains a cycle of length other than $2$. As above, we may remove all multi-edges of $G$ without changing the acyclic orientation number. Moreover, if $G$ contains a loop, we have $t_{G}=0$ by Lemma 2.3. We therefore assume, without loss of generality, that $G$ is simple and contains a cycle of length $k\geq 3$. Let $e$ be an edge in such a cycle. Note that $e$ is not a bridge. Moreover, since $G$ is simple, contracting $e$ yields a graph which is loop-free, while deleting $e$ yields a simple (connected) graph. Therefore, applying the Deletion-contraction Equation (1) at edge $e$, combined with Lemma 2.3, we get $t_{G}=t_{G\setminus e}+t_{G\cdot e}\geq 1+1=2$, as desired. ∎
@@
Another way of viewing this procedure is that it removes “tree branches” that were attached at some vertices in the graph $G$. An example is shown on Figure 1 below. Note that the output $\mathrm{Core}_{2}\left([\right)G]$ only depends on the choice of vertex $v$ in Step (2) if at some stage of the process we reach a graph $H$ which is reduced to a single edge, which occurs in fact exactly when the original graph $G$ is a tree. In that case, removing either of the two vertices will yield a graph consisting of a single vertex, and we simply set $\mathrm{Core}_{2}\left([\right)G]$ to be any arbitrary vertex of the original graph.
$G$$\mathrm{Core}_{2}\left([\right)G]$
@@
For any graph $G$, we have $t_{G}=t_{\mathrm{Core}_{2}\left([\right)G]}$.
@@
Let $G$ be a graph with a vertex $v$ of degree $1$, and consider $e$ the unique edge incident to $v$. By definition, we have $G\setminus\{v\}=G\cdot e$, and $e$ is a bridge in $G$ (deleting it disconnects $v$ from the rest of the graph). Therefore we have $t_{G\setminus\{v\}}=t_{G\cdot e}=t_{G}$ by Equation (1) (bridge case). The result follows by iterating the above until no vertices of degree $1$ remain. ∎
@@
If $G$ is a decorated $k$-cycle for some $k\geq 1$, then we have $t_{G}=k-1$.
@@
If $G$ is a decorated $k$-cycle, then $\mathrm{Core}_{2}\left([\right)G]$ is the $k$-cycle $C_{k}$ by definition. Lemma 2.5 then yields that $t_{G}=t_{C_{k}}$. Applying the Deletion-contraction Equation (1) gives $t_{C_{k}}=t_{P_{k}}+t_{C_{k-1}}$ (where $P_{k}$ is the path graph on $k$ vertices). Since path graphs are trees, Lemma 2.4 implies that $t_{P_{k}}=1$ for any $k\geq 1$, so that $t_{C_{k}}=k-1+t_{C_{1}}$. But $C_{1}$ is the graph consisting of a single loop, so Lemma 2.3 implies that $t_{C_{1}}=0$, and the result follows immediately. ∎
@@
Let $G$ be a simple graph, with a pair of distinct vertices $v,w$ such that $(v,w)$ is not an edge of $G$. Let $G^{\prime}:=G\cup\{(v,w)\}$ be the graph $G$ to which we add the edge $(v,w)$. Then we have $t_{G}<t_{G^{\prime}}$.
@@
Let $G$, $e:=(v,w)$, and $G^{\prime}$ be as in the statement of the lemma, i.e. $G=G^{\prime}\setminus e$. By the Deletion-contraction Equation (1), we have $t_{G^{\prime}}=t_{G}+t_{G^{\prime}\cdot e}$. Moreover, by construction, since $G$ is simple, the graph $G^{\prime}\cdot e$ is loop-free. The result then follows from Lemma 2.3. ∎
@@
Let $G$ be a simple graph, and $G^{\prime}$ a subgraph of $G$. Then we have $t_{G^{\prime}}\leq t_{G}$. Moreover, if the $2$-cores $\mathrm{Core}_{2}\left([\right)G]$ and $\mathrm{Core}_{2}\left([\right)G^{\prime}]$ are not isomorphic, then we have $t_{G^{\prime}}<t_{G}$.
@@
By construction, the above process outputs a spanning subgraph $G^{0}$ of $G$ such that $\mathrm{Core}_{2}\left([\right)G^{0}]$ and $\mathrm{Core}_{2}\left([\right)G^{\prime}]$ are isomorphic (constructing $G^{0}$ from $G^{\prime}$ only adds “tree branches”). In particular by Lemma 2.5, we have $t_{G^{0}}=t_{G^{\prime}}$. Then the graph $G$ can be reached from $G^{0}$ through a (possibly non-empty) sequence of edge additions. Moreover, if $\mathrm{Core}_{2}\left([\right)G]$ and $\mathrm{Core}_{2}\left([\right)G^{\prime}]$ are not isomorphic, there must be at least one edge in $E(G)$ which is not in $E(G^{0})$, in which case this sequence must be non-empty. The results then follow from Lemma 2.7. ∎

### v4

@@ Abstract.
We study a link between complete non-ambiguous trees (CNATs) and permutations exhibited by Daniel Chen and Sebastian Ohlig in recent work. In this, they associate a certain permutation to the leaves of a CNAT, and show that the number of $n$-permutations that are associated with exactly one CNAT is $2^{n-2}$. We connect this to work by the first author and co-authors linking complete non-ambiguous trees and the acyclic orientation number of the associated permutation graph. This allows us to prove a number of conjectures by Chen and Ohlig on the number of $n$-permutations that are associated with exactly $k$ CNATs for various $k>1$, via various bijective correspondences between such permutations. We also exhibit a new bijection between (n1)(n-1)-permutations and CNATs whose permutation is the decreasing permutation $n(n-1)\cdots 1$. This bijection maps the left-to-right minima of the permutation to dots on the top row of the corresponding CNAT, and descents of the permutation to empty rows of the CNAT.
Keywords: complete non-ambiguous trees, permutations, permutation graphs, acyclic orientations.
@@ 1. Introduction
We can associate a permutation to a CNAT by keeping only its leaf dots (see Section 2.4). This link between CNATs and permutations was first noted, to the best of our knowledge, in [13] (see Section 4 in that paper for more details). Chen and Ohlig [11] initiated the first in-depth combinatorial study of this relationship. In particular, they characterised the set of permutations $\pi$ which are associated to a unique CNAT. More generally, they looked into the number of $n$-permutations with $k$ associated CNATs, for $k\geq 1$, and provided a number of conjectures on the enumeration of such permutations. They also studied in detail so-called upper-diagonal CNATs, which are CNATs where the associated permutation is the decreasing permutation $n(n-1)\cdots 1$, using these to prove a conjecture of Laborde-Zubieta on occupied corners of tree-like tableaux [19]. Finally, they introduced a natural statistic, determinant, on CNATs, and conjectured that CNATs with determinant +1+1 and those with determinant 1-1 are equinumerous. This conjecture was solved very recently by Aval [2] through a bijective approach.
@@
In this paper, we will further develop the study initiated in [11], using results from [13]. In that work it was shown that the number of CNATs associated with a given permutation $\pi$ is equal to the number of minimal recurrent configurations for the Abelian sandpile model on the permutation graph $G_{\pi}$ of $\pi$. These are counted by certain acyclic orientations of the graph (see Theorem 2.13 in this paper). This enumerative result allows us to solve a number of conjectures in [11], via various bijective correspondences between permutations with a given number of CNATs.
@@
We also add to the study of upper-diagonal CNATs by providing a new bijection between upper-diagonal CNATs of size $n$ and permutations of length n1n-1 (see Theorem 3.6). This new bijection has the following added benefits. Firstly, it is direct, whereas the bijection in [11] used intermediate objects called tiered trees (see [12]). Secondly, it maps certain statistics of the upper-diagonal CNATs such as the number of top-row dots, or the number of empty rows, to well-known statistics of the corresponding permutation.
@@
Our paper is organised as follows. In Section 2, we recall some necessary definitions of, and notation on, the combinatorial objects that will be considered in this paper. These include graphs, acyclic orientations, permutations, and CNATs with associated permutations. We also show a number of useful preliminary results that will be useful in the remainder of the paper. At the end of the section we provide the key connection between CNATs and the acyclic orientation number of an associated graph (Theorem 2.13). In Section 3, we focus on one specific family of CNATs, called upper-diagonal CNATs. We introduce a concept of labelled CNAT, and describe a bijection between labelled CNATs and permutations of the label set (Theorem 3.6), which preserves certain statistics of both objects. This bijection then specialises to a bijection between upper-diagonal CNATs of size $n$ and permutations of length n1n-1. The bijection relies on two operations on labelled CNATs, called top-row decomposition and top-row deletion.
@@
In Section 4, we focus on counting permutations according to their number of associated CNATs. Specifically, for $k,n\geq 1$ we are interested in the set $B(n,k)$ of permutations of length $n$ which are associated with exactly $k$ CNATs. The study of these sets was initiated in [11], and we continue that work here by proving a number of conjectures left by the authors. We begin in Section 4.1 by giving characterisations of the sets $B(n,1)$ and $B(n,2)$ in terms of the so-called quadrant condition (Definition 4.1, Propositions 4.2 and 4.9). This allows us to describe a simple bijection between the product set $\{2,\dots,n\}\times B(n,1)$ and $B(n+1,2)$ (Theorem 4.3), and deduce the enumerative formula for the latter (Corollary 4.5). In Section 4.2 we use permutation patterns to establish a bijection between the sets $B(n,2)$ and $B(n+1,3)$ (Theorem 4.15). We then show in Section 4.3 that the set $B(n,5)$ is always empty (Theorem 4.17). Finally, in Section 4.4 we consider the maximal value of $k$ such that $B(n,k)$ is non-empty. We show (Theorem 4.23) that this is achieved for k=(n1)!k=(n-1)!, and that the unique permutation in B(n,(n1)!)B(n,(n-1)!) is precisely the decreasing permutation $n(n-1)\cdots 1$ whose CNATs were studied in detail in Section 3. We conclude our paper in Section 5 with a brief summary of our results as well as some open problems and directions for future research.
@@ 2.1. Graphs and orientations
A cycle in the graph $G$ is a sequence of vertices $v_{0},v_{1},\ldots,v_{n}=v_{0}$ (for some $n\geq 3$) such that for every $i>0$, $(v_{i-1},v_{i})$ is an edge of $G$, and the vertices $v_{0},v_{1},\ldots,v_{n-1}$ are all distinct. The integer $n$ is the length of the cycle, and we will say that $G$ contains a cycle of length $n$ (or $n$-cycle for short) if there exists such a cycle in $G$. The $n$-cycle (graph) $C_{n}$ is the graph consisting of a single cycle of length $n$, with no other vertices or edges. If $G$ contains a cycle $v_{0},v_{1},\ldots,v_{n}=v_{0}$ such that $G\left[\{v_{0},\ldots,v_{n-1}\}\right]$ is the $n$-cycle $C_{n}$, we say that $G$ induces an $n$-cycle. A tree is a (connected) graph containing no cycle of any length. A spanning tree of a graph $G$ is a spanning subgraph which is a tree. The $n$-path (graph) is the graph with vertex set $V(G)=\{v_{1},\ldots,v_{n}\}$ and edge set $E(G)=\{(v_{i},v_{i+1})\}_{1\leq i\leq n-1}$ (equivalently the $n$-cycle with any single edge removed).
@@
An orientation is acyclic if it contains no directed cycle, i.e. there is no sequence $v_{0},\ldots,v_{n-1}$ of vertices such that $v_{0}\xrightarrow{\mathcal{O}}v_{1}\xrightarrow{\mathcal{O}}\cdots\xrightarrow{\mathcal{O}}v_{n-1}\xrightarrow{\mathcal{O}}v_{0}$. It is straightforward to check that an acyclic orientation must have at least one source and at least one sink. For $s\in V(G)$, we see that an acyclic orientation $\mathcal{O}$ is $s$-rooted if $s$ is the unique sink of $\mathcal{O}$.
2.2. Acyclic orientation numbers
A key result that we will use in this paper links CNATs (see Section 2.4) to a certain quantity related to acyclic orientations of graphs.
@@ 2.2. Acyclic orientation numbers
Let $G$ be a graph, and $s\in V(G)$ a fixed sink vertex of $G$. Then the number of $s$-rooted acyclic orientations of $G$ does not depend on the choice of sink vertex $s$. We denote this number $a_{G}$, and call it the acyclic orientation number.
Proof.
Let $T_{G}$ denote the Tutte polynomial of a graph $G$ (see e.g. [14, Chapter 2] for definitions and more information on this important polynomial). Then we have that $T_{G}(1,0)$ is equal to the number of $s$-rooted acyclic orientations of $G$ for any fixed $s\in V(G)$ (see e.g. [6] for a bijective proof of this fact). Thus $a_{G}$ does not depend on the choice of sink vertex $s$. ∎
We first recall the famous deletion-contraction relation for the acyclic orientation number of a graph $G$. This is a simple specialisation of the more general result for the Tutte polynomial (see e.g. [14, Definition 2.7]). Given a graph $G$ and an edge $e=(v,w)\in E(G)$, we denote by $G\setminus e$ the graph $G$ with edge $e$ deleted. That is, $G\setminus e:=\left(V(G),E(G)\setminus\{e\}\right)$. Similarly, $G\cdot e$ denotes the graph $G$ with edge $e$ contracted. This is the graph $G$ in which the two end-points $v$ and $w$ are replaced by a single vertex $vw$, and edges incident to either $v$ or $w$ are made incident to $vw$ instead (with multiplicity preserved).
Proposition 2.2.
Given a graph $G$ and an edge $e\in E(G)$, the acyclic orientation number $a_{G}$ satisfies the following recurrence:
[…]
We now state a number of useful structural results on how the acyclic number of a graph behaves with respect to certain graph constructions. We begin by characterising graphs whose acyclic orientation number is 0 or $1$.
@@
Let $G$ be a graph. Then we have $a_{G}=0$ if, and only if, $G$ contains a loop. In particular, if $G$ is simple, we have $a_{G}\geq 1$.
@@
If $G$ contains a loop, then Equation (1) implies that $a_{G}=0$. Conversely, if $G$ is loop-free, fix some arbitrary total order $v_{1},\ldots,v_{n}$ of the vertex set $V(G)$. Define an orientation $\mathcal{O}$ of $G$ by $v_{i}\xrightarrow{\mathcal{O}}v_{j}$ if, and only if, $i<j$. Since $G$ is loop-free, this does indeed assign a direction to each edge of $G$, and by construction $\mathcal{O}$ is a $v_{n}$-rooted acyclic orientation. Proposition 2.1 then implies that $a_{G}\geq 1$, as desired. ∎
@@
Let $G$ be a graph. Then we have $a_{G}=1$ if, and only if, all cycles in $G$ have length $2$. In particular, if $G$ is simple, then $a_{G}=1$ if, and only if, $G$ is a tree.
@@
Suppose that $G$ is such that all its cycles have length $2$, and let $e=(v,w)$ be a multi-edge corresponding to such a cycle (i.e. $e$ has multiplicity at least $2$ in $G$). By definition, we have $a_{G}=a_{G\setminus e}+a_{G\cdot e}$. But contracting $e$ in $G$ yields a loop at the new vertex $vw$, since $e$ was a multi-edge in $G$. Therefore $a_{G\cdot e}=0$ by Lemma 2.3, which implies that $a_{G}=a_{G\setminus e}$. Repeating this process until all mutli-edges have been removed, we are left with a tree $G^{\prime}$ and $a_{G}=a_{G^{\prime}}$. But now all edges in $G^{\prime}$ are bridges, and we may contract them one-by-one until we reach a graph consisting of a single vertex. Applying Equation (1) throughout this process immediately yields $a_{G^{\prime}}=1$, as desired.
Conversely, suppose that $G$ contains a cycle of length other than $2$. As above, we may remove all multi-edges of $G$ without changing the acyclic orientation number. Moreover, if $G$ contains a loop, we have $a_{G}=0$ by Lemma 2.3. We therefore assume, without loss of generality, that $G$ is simple and contains a cycle of length $k\geq 3$. Let $e$ be an edge in such a cycle. Note that $e$ is not a bridge. Moreover, since $G$ is simple, contracting $e$ yields a graph which is loop-free, while deleting $e$ yields a simple (connected) graph. Therefore, applying the Deletion-contraction Equation (1) at edge $e$, combined with Lemma 2.3, we get $a_{G}=a_{G\setminus e}+a_{G\cdot e}\geq 1+1=2$, as desired. ∎
@@
Another way of viewing this procedure is that it removes “tree branches” that were attached at some vertices in the graph $G$. An example is shown in Figure 1 below. Note that the output $\mathrm{Core}_{2}\left([\right)G]$ only depends on the choice of vertex $v$ in Step (2) if at some stage of the process we reach a graph $H$ which is reduced to a single edge, which occurs in fact exactly when the original graph $G$ is a tree. In that case, removing either of the two vertices will yield a graph consisting of a single vertex, and we simply set $\mathrm{Core}_{2}\left([\right)G]$ to be any arbitrary vertex of the original graph.
$G]\mathrm{Core}_{2}\left([\right)G$]
@@
For any graph $G$, we have $a_{G}=a_{\mathrm{Core}_{2}\left([\right)G]}$.
@@
Let $G$ be a graph with a vertex $v$ of degree $1$, and consider $e$ the unique edge incident to $v$. By definition, we have $G\setminus\{v\}=G\cdot e$, and $e$ is a bridge in $G$ (deleting it disconnects $v$ from the rest of the graph). Therefore we have $a_{G\setminus\{v\}}=a_{G\cdot e}=a_{G}$ by Equation (1) (bridge case). The result follows by iterating the above until no vertices of degree $1$ remain. ∎
@@
If $G$ is a decorated $k$-cycle for some $k\geq 1$, then we have $a_{G}=k-1$.
@@
If $G$ is a decorated $k$-cycle, then $\mathrm{Core}_{2}\left([\right)G]$ is the $k$-cycle $C_{k}$ by definition. Lemma 2.5 then yields that $a_{G}=a_{C_{k}}$. Applying the Deletion-contraction Equation (1) gives $a_{C_{k}}=a_{P_{k}}+a_{C_{k-1}}$ (where $P_{k}$ is the path graph on $k$ vertices). Since path graphs are trees, Lemma 2.4 implies that $a_{P_{k}}=1$ for any $k\geq 1$, so that $a_{C_{k}}=k-1+a_{C_{1}}$. But $C_{1}$ is the graph consisting of a single loop, so Lemma 2.3 implies that $a_{C_{1}}=0$, and the result follows immediately. ∎
@@
Let $G$ be a simple graph, with a pair of distinct vertices $v,w$ such that $(v,w)$ is not an edge of $G$. Let $G^{\prime}:=G\cup\{(v,w)\}$ be the graph $G$ to which we add the edge $(v,w)$. Then we have $a_{G}<a_{G^{\prime}}$.
@@
Let $G$, $e:=(v,w)$, and $G^{\prime}$ be as in the statement of the lemma, i.e. $G=G^{\prime}\setminus e$. By the Deletion-contraction Equation (1), we have $a_{G^{\prime}}=a_{G}+a_{G^{\prime}\cdot e}$. Moreover, by construction, since $G$ is simple, the graph $G^{\prime}\cdot e$ is loop-free. The result then follows from Lemma 2.3. ∎
@@
Let $G$ be a simple graph, and $G^{\prime}$ a subgraph of $G$. Then we have $a_{G^{\prime}}\leq a_{G}$. Moreover, if the $2$-cores $\mathrm{Core}_{2}\left([\right)G]$ and $\mathrm{Core}_{2}\left([\right)G^{\prime}]$ are not isomorphic, then we have $a_{G^{\prime}}<a_{G}$.
@@
By construction, the above process outputs a spanning subgraph $G^{0}$ of $G$ such that $\mathrm{Core}_{2}\left([\right)G^{0}]$ and $\mathrm{Core}_{2}\left([\right)G^{\prime}]$ are isomorphic (constructing $G^{0}$ from $G^{\prime}$ only adds “tree branches”). In particular by Lemma 2.5, we have $a_{G^{0}}=a_{G^{\prime}}$. Then the graph $G$ can be reached from $G^{0}$ through a (possibly empty) sequence of edge additions. Moreover, if $\mathrm{Core}_{2}\left([\right)G]$ and $\mathrm{Core}_{2}\left([\right)G^{\prime}]$ are not isomorphic, there must be at least one edge in $E(G)$ which is not in $E(G^{0})$ (otherwise we would have $G^{0}=G$ and therefore $\mathrm{Core}_{2}\left([\right)G]=\mathrm{Core}_{2}\left([\right)G^{0}]$ and $\mathrm{Core}_{2}\left([\right)G^{\prime}]$ would be isomorphic), in which case this sequence must be non-empty. The results then follow from Lemma 2.7. ∎